OpenStudy (anonymous):

Math please help! What is the value of d + e + f when d = 20, e = –4, and f = –2? A. –26 B. –22 C. 14 D. 18

OpenStudy (anonymous):

20-4-2=20-6=14 option C...!! @keepsecure93

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

yw